what is equivalent fractions for 6/5?how do you get the answer
Answer #1
u divide 6 by 5 or multiply both numbers in the fraction by the same number
answered by: starzy
Answer #2
multiplying the top and bottom of 6/5 by the same number will result in a new equivalent fraction.

e.g.
6/5 = 12/10 = 18/15 = 60/50 = ...
answered by: Tinu
Answer #3
6/5=12/10=18/15=60/50
source: Maths
answered by: Vaibhav choudhary
Answer #4

12/10 and 18/15
Answer #5
6/5=12/10=18/15=60/50
source: My brain
answered by: Joyce